# ReaGAN: Node-as-Agent-Reasoning Graph Agentic Network

### v2

@@ 3 Experiments
  • RQ3: When do agentic nodes need global semantic retrieval to do global aggregation, and how should local and global shots be balanced in prompts?
@@
Table 1: Test accuracy (%) on the Cora, Citeseer, and Chameleon datasets. Traditional GNNs rely on parametric training and fixed message passing, while ReaGAN leverages a frozen LLM for agentic planning and personalized reasoning.
@@
We compare ReaGAN with a set of widely used baselines, including GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), GraphSAGE (Hamilton et al., 2017), APPNP (Klicpera et al., 2019), GPRGNN (Chien et al., 2021), MixHop (Abu-El-Haija et al., 2019), and MLP-2. All models are implemented using PyTorch Geometric with standard hyperparameters and are trained on the same data splits as ReaGAN.
@@
We conduct all experiments in PyTorch using a cluster of NVIDIA RTX A6000 GPUs. The agentic node reasoning module relies on a frozen LLM, served via vLLM, using Qwen2-14B. We do not perform any fine-tuning. We report test accuracy for node classification. To ensure fair comparison, all baselines are trained under identical splits. We use all-MiniLM-L6-v2 to transfer text to embedding. More experimental details are provided in Appendix.
3.1 Overall Performance on Node Classification
To answer RQ1, we report node classification accuracy for all baselines and for ReaGAN in Table 1. Despite using no trainable parameters, ReaGAN achieves competitive performance compared to fully supervised GNNs, while enabling node-level autonomy and reasoning via a frozen LLM.
3.2 Impact of Agentic Planning and Global Retrievals
To answer RQ2, we perform ablations on ReaGAN’s two core components: node-level planning and global retrieval. Removing prompt planning forces all nodes to follow a fixed action sequence, leading to notable accuracy drops due to the lack of context-specific behavior. Disabling global retrieval (Local Only) limits each node to structural neighbors, which underperforms in sparse graphs. Conversely, the Global Only variant removes structural input and struggles in well-connected graphs, showing the need for both structural and semantic signals. Experimental results are shown on Table 2 which shows that both agentic planning and global semantic access are essential. Removing either component leads to measurable accuracy degradation, validating the effectiveness of our solution to the autonomy and semantic challenges.
Table 2: Ablation study on Cora, Citeseer, and Chameleon (Test Accuracy %). ReaGAN’s performance depends on both agentic planning and local-global integration.
@@ 3.2 Impact of Agentic Planning and Global Retrievals
Table 3: Prompt Memory Strategy vs. Accuracy. Comparison of two prompt construction strategies across three datasets. Strategy A includes both local and global memory in the prompt. Strategy B includes global memory only when fewer than two local entries are available.
@@
Chameleon B 38.29
To address RQ3, we investigate how the balance between local and global memory in the prompt affects node classification accuracy. We perform an ablation study comparing two prompt construction strategies: Strategy A includes both local and global memory in the prompt; Strategy B includes global memory only when fewer than two local examples are available. As shown in Table 3, Strategy A consistently performs better on structurally dense graphs like Cora and Chameleon, where high-quality local memory is readily available. In contrast, on the sparser Citeseer graph, Strategy B outperforms Strategy A by selectively avoiding potentially noisy global memory when local context is sufficient. The performance gap across datasets can be attributed to differences in structural density and label locality. Cora and Chameleon have stronger local homophily, making global memory broadly useful, while Citeseer benefits more from selective use due to sparser connections and noisier neighborhoods. As for RQ4, Table 4 reveals that exposing label semantics consistently harms accuracy, as LLMs tend to overfit to general label names like “machine learning” and make biased guesses.
Table 4: Effect of Label Semantics on Accuracy. Showing label names (e.g., "Rule Learning") in the prompt harms performance, as LLMs tend to overfit to label wording rather than reasoning from memory. We anonymize all labels (e.g., "Label_2") in our final design.
@@
4 Related Works
4.1 Graph Neural Networks and Message Passing
@@
Graph Neural Networks (GNNs) such as GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), and GAT (Veličković et al., 2018) have become the dominant paradigm in graph machine learning. These models rely on fixed, layer-wise message passing schemes that aggregate information from each node’s neighbors using predefined update functions. While effective, this rigid design limits expressiveness and can cause issues like over-smoothing and over-squashing. To overcome these limitations, several works have proposed more flexible message passing strategies. CoGNN (Finkelshtein et al., 2023) introduces cooperative agents that decide whether to broadcast or listen during each round, enabling more adaptive communication. However, these agents still rely on hand-crafted utility functions and rule-based execution. In contrast, our method allows each node to independently plan and execute its own message passing actions using an LLM-based agentic mechanism, and extends the communication space to include global semantic neighbors, not just local structural ones.
4.2 Large Language Models for Graph Tasks
@@
Recent work has explored the use of LLMs for graph learning (Shu et al., 2024; Zhao et al., 2024). PromptGFM (Zhu et al., 2024) converts nodes into text-based prompts and uses LLMs to learn a graph vocabulary for classification. In-context RAG (Li et al., 2024) frames node classification as a retrieval-augmented generation task, where textual neighbors are fetched and provided to the LLM for reasoning. These methods effectively leverage LLM capabilities but do not enable node-level autonomy or action planning. By contrast, we use LLMs not as passive inference engines but as active planners—the cognitive core of each agent node. The LLM decides what action to take next, what context to gather, and when to predict.
4.3 Agent-Based Graph Learning
@@
Several prior studies have introduced the notion of “agents” in graphs, but differ from our formulation. AgentNet (Martinkus et al., 2023) trains neural agents to walk the graph and distinguish structures via learned exploration policies, but these agents are part of a supervised model and do not make autonomous decisions. GraphAgent (Zhang et al., 2023) and GAgN (Liu et al., 2024) also propose node-agent architectures for adversarial defense and resilience, but use hardcoded 1-hop views or restricted inference logic rather than learned behaviors. AgentGNN (Liu et al., 2022) applies the agent-based paradigm to spatiotemporal graphs, modeling each node as a temporal processor. While similar in spirit, their approach is limited to sequential data and does not support open-ended, action-driven interaction or reasoning. Our formulation is fundamentally different: we treat each node as a full-fledged intelligent agent, with the ability to observe, reason, and act using LLM-powered prompts, without predefined roles, hardcoded transitions, or limited interaction space.
5 Conclusion
@@
6 Limitations and Future Work
@@
In future work, we envision ReaGAN as a foundational layer in multi-agent systems (Li et al., 2025), where each node-agent can participate in decentralized decision-making(Huang et al., 2025) or inter-agent communication. For instance, its ability to reason locally and retrieve information globally makes it a strong fit for modular systems (Gershon et al., 2025; Jin et al., 2025) or routing-based architectures (Mei et al., 2025a), where nodes may act as autonomous modules coordinating across a shared substrate. Additionally, to scale ReaGAN under constrained resources, we propose integrating it with agent orchestration frameworks such as AIOS (Mei et al., 2025b)—enabling parallelized, resource-aware scheduling of reasoning agents across large graphs. Rather than being limited to graph classification, we see ReaGAN as a general blueprint for scalable agentic inference—well suited for future systems that demand fine-grained, context-aware, and communication-capable reasoning units.
@@ Appendix A Technical Appendices and Supplementary Material
A.1 Prompt Example
A.1.1 Planning Prompt Example
@@
  {"action_type": "local aggregate", "global aggregate" or "no_op"},
  {"action_type": "local aggregate", "global aggregate" or "no_op"}
A.1.2 Prediction Prompt Example
You are a label prediction agent.
You will be given a new node’s aggregated text feature along with a memory of labeled examples. Use the memory to infer the most likely label for this node. Respond strictly in the required JSON format.
Node State: - Text Feature: “Adverse interaction with tree depth restriction.”
@@ A.1.2 Prediction Prompt Example
1. Analyze the Current Node Text: - Identify primary topics and application domain - Determine the specific problem being solved - Note core methodologies and algorithms"
@@ 4. Avoid over-reliance on isolated keywords
First analyze memory examples and their labels, then compare them to the input text. Identify the most semantically similar memory items and explain why. Finally, decide which label best matches and explain your reasoning.
@@
{"action_type": "predict", "predicted_label": "Label ID"}
A.2 Dataset Construction and Processing
@@
Table 5: Dataset statistics for node classification benchmarks.
@@
A.3 Hyperparameter Settings
ReaGAN.

### v4

@@ 3. Experiments
  • RQ3: When do agentic nodes need global semantic retrieval to do global aggregation, and how should local and global examples be balanced in prompts?
@@
Table 1. Test accuracy (%) on the Cora, Citeseer, and Chameleon datasets. Traditional GNNs rely on parametric training and fixed message passing, while ReaGAN leverages a frozen LLM for agentic planning and personalized reasoning.
@@
We compare ReaGAN with a set of widely used baselines, including GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), GraphSAGE (Hamilton et al., 2017), APPNP (Klicpera et al., 2019), GPRGNN (Chien et al., 2020), MixHop (Abu-El-Haija et al., 2019), and MLP-2. All models are implemented using PyTorch Geometric with standard hyperparameters and are trained on the same data splits as ReaGAN.
@@
We conduct all experiments in PyTorch using a cluster of NVIDIA RTX A6000 GPUs. The agentic node reasoning module relies on a frozen LLM, served via vLLM, using Qwen2.5-14B-Instruct. We do not perform any fine-tuning. We report test accuracy for node classification. To ensure fair comparison, all baselines are trained under identical splits. We use all-MiniLM-L6-v2 to transfer text to embedding. More experimental details are provided in the Appendix. A.3.
3.1. Overall Performance on Node Classification
To answer RQ1, we compare ReaGAN’s node classification accuracy against a suite of fully supervised GNN baselines, with the results detailed in Table 1. The findings demonstrate that ReaGAN, despite operating with a frozen LLM and without any gradient-based training, achieves remarkably competitive performance. On the Cora dataset, ReaGAN achieves an accuracy of $84.95\%$, slightly outperforming well-established models like GCN ($84.71\%$) and GraphSAGE ($84.35\%$). This result is particularly noteworthy as it shows that an agentic, reasoning-based approach can match or even exceed the performance of models that have been explicitly trained on the graph’s structure and features for this specific task. On the Citeseer and Chameleon datasets, while some supervised methods like GraphSAGE and GPRGNN show higher accuracy, ReaGAN still maintains a strong competitive standing. For instance, its performance on Citeseer ($60.25\%$) is comparable to several established GNNs. It is crucial to frame this result in the context of our method’s paradigm: ReaGAN is not fine-tuned on the downstream task. Instead, it leverages the pre-existing reasoning capabilities of an LLM to interpret the graph context in a zero-shot, few-shot manner. In summary, these results validate the core premise of our work. By empowering nodes with autonomy and integrating both local and global information through a reasoning-based framework, ReaGAN establishes itself as a viable and powerful new paradigm for graph machine learning, one that trades task-specific training for greater flexibility and emergent reasoning capabilities.
3.2. Impact of Agentic Planning and Global Retrievals
To answer RQ2, we perform ablations on ReaGAN’s two core components: node-level planning and global retrieval. Removing prompt planning forces all nodes to follow a fixed action sequence, leading to notable accuracy drops due to the lack of context-specific behavior. Disabling global retrieval (Local Only) limits each node to structural neighbors, which underperforms in sparse graphs. Conversely, the Global Only variant removes structural input and struggles in well-connected graphs, showing the need for both structural and semantic signals. Experimental results are shown on Table 2 which shows that both agentic planning and global semantic access are essential. Removing either component leads to a measurable degradation in accuracy, validating the effectiveness of our solution in addressing autonomy and semantic challenges.
Table 2. Ablation study on Cora, Citeseer, and Chameleon (Test Accuracy %). ReaGAN’s performance depends on both agentic planning and local-global integration.
@@ 3.2. Impact of Agentic Planning and Global Retrievals
Table 3. Prompt Memory Strategy vs. Accuracy. Comparison of two prompt construction strategies across three datasets. Strategy A includes both local and global memory in the prompt. Strategy B includes global memory only when fewer than two local entries are available.
@@
B 38.29
To address RQ3, we investigate how the balance between local and global memory in the prompt affects node classification accuracy. We perform an ablation study comparing two prompt construction strategies: Strategy A, which includes both local and global memory in the prompt, and Strategy B, which includes global memory only when fewer than two local examples are available. The results in Table 3, particularly Citeseer’s contrasting performance, highlight the critical need for an adaptive retrieval strategy. The superior performance of Strategy B on Citeseer is not arbitrary; it is a direct consequence of the underlying graph topology. A quantitative analysis reveals that Citeseer is a significantly sparser graph than Cora. Based on the dataset statistics in Table 5, Citeseer has an average node degree of approximately 2.84 ($4,732\text{ edges}/3,327\text{ nodes}\times 2$), which is considerably lower than Cora’s average degree of roughly 4.01 ($5,429\text{ edges}/2,708\text{ nodes}\times 2$). This structural sparsity has a crucial implication: many nodes in Citeseer have small, less informative local neighborhoods. In such a setting, Strategy A—which indiscriminately performs global retrieval—risks introducing “semantic noise.” While retrieved nodes may be textually similar, they might lack the precise contextual relevance that structural proximity provides, thereby confusing the agent and degrading performance. In contrast, Strategy B functions as a more intelligent and cautious mechanism. It prioritizes the high-fidelity signals from the immediate structural neighborhood and only resorts to global retrieval as a fallback when local information is critically insufficient (i.e., fewer than two local examples are available). This selective approach prevents the agent from diluting a sufficient local context with potentially noisy global information, leading to a more robust and accurate classification on sparser, “noisier” graphs like Citeseer.
Regarding RQ4, we investigate the impact of label semantics on the agent’s reasoning process. The results, presented in Table 4, are unequivocal: providing explicit, human-readable label names (e.g., ”Machine Learning”, ”Databases”) consistently degrades classification accuracy across all datasets. This performance drop reveals a critical insight into the behavior of LLMs in graph-based reasoning tasks. When exposed to semantic label names, the LLM leverages its vast pre-trained knowledge to take a cognitive “shortcut.” Instead of reasoning from the evidence provided in the prompt—such as the node’s features and the contextual examples from its neighbors—the model tends to overfit to the literal meaning of the labels. For instance, if a node’s text contains keywords like “algorithm” or “prediction,” the LLM might impulsively select the “Machine Learning” label, even if the local and global examples in its memory point towards a more specific or different category. This creates biased guesses based on superficial keyword matching rather than a holistic analysis of the provided context. Conversely, by anonymizing the labels (e.g., “Label_1”, “Label_2”), we compel the agent to perform a more genuine and robust reasoning task. Without the semantic shortcut, the LLM is forced to rely entirely on the in-context examples provided in the prompt to understand the defining characteristics of each category. It must compare the target node’s features against the features of the labeled neighbors to infer the correct classification. This process ensures that the model’s predictions are grounded in the specific, contextual evidence from the graph structure and node attributes, rather than its pre-existing and often overly general world knowledge. Our final design therefore adopts label anonymization as a crucial step to ensure faithful reasoning.
Table 4. Effect of Label Semantics on Accuracy. Showing label names (e.g., ”Rule Learning”) in the prompt harms performance, as LLMs tend to overfit to label wording rather than reasoning from memory. We anonymize all labels (e.g., ”Label_2”) in our final design. we only reveal label names in RQ4 for this controlled comparison.
@@
4. Related Work
4.1. Graph Neural Networks and Message Passing
@@
Graph Neural Networks (GNNs) such as GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), and GAT (Veličković et al., 2018) have become the dominant paradigm in graph machine learning. These models rely on fixed, layer-wise message passing schemes that aggregate information from each node’s neighbors using predefined update functions. While effective, this rigid design limits expressiveness and can cause issues like over-smoothing and over-squashing. To overcome these limitations, several works have proposed more flexible message passing strategies. CoGNN (Finkelshtein et al., 2023) introduces cooperative agents that decide whether to broadcast or listen during each round, enabling more adaptive communication. However, these agents still rely on handcrafted utility functions and rule-based execution. In contrast, our method allows each node to independently plan and execute its own message passing actions sing an LLM-based agentic mechanism, and extends the communication space beyond local structural neighbors to include global semantic neighbors.
4.2. Large Language Models for Graph Tasks
@@
Recent work has explored the use of LLMs for graph learning (Shu et al., 2024; Zhao et al., 2024). PromptGFM (Zhu et al., 2024) converts nodes into text-based prompts and uses LLMs to learn a graph vocabulary for classification. In-context RAG (Li et al., 2024) frames node classification as a retrieval-augmented generation task, where textual neighbors are fetched and provided to the LLM for reasoning. These methods effectively leverage LLM capabilities but do not enable node-level autonomy or action planning. STGP (Hu et al., 2024), a prompt-based transfer learning framework for spatio-temporal graphs, to tackle the challenge of cross-domain and cross-task generalization. By using a unified task template and a two-stage prompting mechanism, the framework allows a model to effectively adapt to multiple tasks in new, data-scarce domains, achieving significant performance improvements on tasks such as forecasting and kriging. By contrast, we use LLMs not as passive inference engines but as active planners—the cognitive core of each agent node. The LLM decides what action to take next, what context to gather, and when to predict.
4.3. Agent-Based Graph Learning
@@
Several prior studies have introduced the notion of “agents” in graphs, but differ from our formulation. AgentNet (Martinkus et al., 2023) trains neural agents to walk the graph and distinguish structures via learned exploration policies, but these agents are part of a supervised model and do not make autonomous decisions. GAgN (Liu et al., 2024) also proposes node-agent architectures for adversarial defense and resilience, but use hardcoded 1-hop views or restricted inference logic rather than learned behaviors. Our formulation is fundamentally different: we treat each node as a full-fledged intelligent agent, with the ability to observe, reason, and act using LLM-powered prompts, without predefined roles, hardcoded transitions, or limited interaction space.
5. Conclusion
@@
6. Limitations and Future Work
@@
In future work, we envision ReaGAN as a foundational layer in multi-agent systems (Li et al., 2025b; Chen et al., 2025), where each node-agent can participate in decentralized decision-making(Huang et al., 2025) or inter-agent communication. For instance, its ability to reason locally and retrieve information globally makes it a strong fit for modular systems (Gershon et al., 2025) or routing-based architectures (Mei et al., 2025a), where nodes may act as autonomous modules coordinating across a shared substrate. Additionally, to scale ReaGAN under constrained resources, we propose integrating it with agent orchestration frameworks such as AIOS (Mei et al., 2025b)—enabling parallelized, resource-aware scheduling of reasoning agents across large graphs. Rather than being limited to graph classification, we see ReaGAN as a general blueprint for scalable agentic inference—well suited for future systems that demand fine-grained, context-aware, and communication-capable reasoning units.
7. Acknowledgment
We would like to thank Mingyu Jin, Kai Mei, Wujiang Xu, and Ziwei Zhao for their insightful discussions and valuable feedback, which greatly contributed to the refinement of our framework.
@@ Appendix A Technical Appendices and Supplementary Material
A.1. Prompt Example
A.1.1. Planning Prompt Example
@@
  {"action_type": "local aggregate",
   "global aggregate" or "no_op"},
  {"action_type": "local aggregate",
    "global aggregate" or "no_op"}
A.1.2. Prediction Prompt Example
You are a label prediction agent. You will be given a new node’s aggregated text feature along with a memory of labeled examples. Use the memory to infer the most likely label for this node. Respond strictly in the required JSON format.
Node State:
- Text Feature: “Adverse interaction with tree depth restriction.”
@@ A.1.2. Prediction Prompt Example
1. Analyze the Current Node Text: - Identify primary topics and application domain - Determine the specific problem being solved - Note core methodologies and algorithms”
@@ 4. Avoid over-reliance on isolated keywords
First, analyze memory examples and their labels, then compare them to the input text. Identify the most semantically similar memory items and explain why. Finally, decide which label best matches and explain your reasoning.
@@
{"action_type": "predict",
 "predicted_label": "Label ID"}
A.2. Dataset Construction and Processing
@@
Table 5. Dataset statistics for node classification benchmarks.
@@
A.3. Hyperparameter Settings
ReaGAN Settings